# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ 4.3 Latent and Pixel Space Model: CelebA and ImageNet
[…]
We evaluate the inpainting performance of LanPaint on the CelebA-HQ-256 Liu et al. (2015) and ImageNet-256 Deng et al. (2009) datasets, leveraging pre-trained latent Rombach et al. (2021) and pixel space Dhariwal & Nichol (2021) diffusion models, respectively. The experiments assess reconstruction quality across various mask geometries, including box, half, checkerboard, and outpainting. Following the setting of the previous works (Kawar et al., 2022; Chung et al., 2022b), perceptual fidelity is quantified through LPIPS Zhang et al. (2018) and FID metrics, calculated on 1,000 validation images per dataset, with results presented in Tables 1 and 2. We also provide quanlititive visualization of generated samples in Fig.5 and Fig.6. All methods employ consistent parameters across tasks and masks, utilizing a 20-step Euler Discrete Sampler. Further details are provided in Appendix G.
LanPaint consistently achieves superior LPIPS and FID scores across most test scenarios, demonstrating robustness, particularly in challenging checkerboard and outpainting tasks. In contrast, methods such as DPS Chung et al. (2022a) and CoPaint Zhang et al. (2023a), designed for stochastic sampling with 250–500 steps, exhibit reduced performance in the 20-step ODE setting.
Notably, removing the manifold constraint from MCG Chung et al. (2022b) in DPS and CoPaint compromises stability, leading to poorer performance compared to MCG, which remains robust among heuristic methods. This contradicts prior findings in stochastic DDPM sampling, where removing the manifold constraint typically enhances performance.
On CelebA, Replace Song & Ermon (2019b) marginally outperforms LanPaint in FID for the half-mask scenario (30.3 vs. 30.4), a result attributed to FID’s high variance when large inpainted regions deviate significantly from the original images. This variance is exacerbated by the use of 1,000 validation images, as opposed to the typical 30,000 for highly divergent sets, similarly affecting outpainting results. Consequently, FID scores for half-mask and outpainting scenarios should be interpreted cautiously.
Table 2: LPIPS and FID comparison on ImageNet for various inpainting and outpainting setups. Euler Discrete Sampler, 20 steps. Lower LPIPS and FID values indicate better perceptual similarity and feature distribution similarity to the ground truth, respectively. Numerical suffixes (-5, -10) denote inner iterations (network evaluations per sampling step, except for CoPaint which requires multiple evaluations per inner iteration). Time per sample and memory overhead (extra memory required during inference) are also reported. Evaluations were conducted on a single RTX 3090.
@@
LanPaint-5 0.180 49.3 0.323 49.3 0.127 24.1 0.508 68.6 11.3 599
LanPaint-10 0.171 46.4 0.314 44.7 0.117 21.2 0.486 62.0 20.8 599
[…]
@@ 4.4 Ablation Study
Table 3 presents the ablation study of LanPaint’s major components: the Bidirectional Generative (BiG) score and Fast Langevin Dynamics (FLD), along with the impact of step size on performance. The study is conducted on two datasets, CelebA-HQ-256 and ImageNet, using box inpainting tasks. Results are reported for three different step sizes (0.05, 0.1, and 0.15) with LPIPS and FID metrics, where lower values indicate better performance.
For the ImageNet dataset, the ablation study demonstrates that both the BiG score and FLD significantly contribute to overall performance. Without FLD, the Langevin dynamics diverge as the step size increases from 0.05 to 0.15, resulting in progressively worse performance metrics. However, incorporating FLD suppresses this divergence, enabling the use of larger step sizes and improving performance. The BiG score enhances performance by facilitating bidirectional information flow between the inpainted and known regions, while FLD supports larger step sizes, accelerating the convergence of the Langevin dynamics and yielding better results with the same number of iteration steps.
@@
Table 3: Ablation study of LanPaint-10’s components on CelebA-HQ-256 and ImageNet with box inpainting. Results for different step sizes (0.05, 0.1, 0.15) are shown with LPIPS and FID metrics. Lower values are better.
Step Size 0.05 Step Size 0.1 Step Size 0.15
Method LPIPS FID LPIPS FID LPIPS FID
@@
None (Langevin) 0.115 28.4 0.111 29.2 0.114 30.9
+ BiG score 0.104 26.5 0.102 28.5 0.103 28.7
+ FLD 0.115 28.9 0.112 29.9 0.116 31.7
+ (BiG score + FLD) 0.105 26.6 0.103 28.5 0.103 29.5
@@
None (Langevin) 0.223 65.1 0.314 81.6 0.441 125.9
+ BiG score 0.213 58.0 0.303 76.0 0.431 121.4
+ FLD 0.205 60.7 0.195 56.9 0.188 54.4
+ (BiG score + FLD) 0.190 52.7 0.179 48.1 0.171 46.4
@@ 4.5 Production-Level Model Evaluation Across Architectures: Stable Diffusion, Flux, and HiDream
Previous evaluations of LanPaint were limited to academic benchmarks, leaving its performance on real-world production diffusion models—with their diverse architectures and higher resolutions—unexamined. Notably, to the best of our knowledge, no prior training-free inpainting methods, except variants of the replace method (build in ComfyUI), have demonstrated such validation in their publications or been implemented by third parties for this purpose.
@@
To assess LanPaint’s effectiveness on modern generative models, we implemented it on HiDream-L1 (HiDream.ai, 2025), Flux.1 Dev (Labs, 2024), Stable Diffusion 3.5 (Esser et al., 2024), and Stable Diffusion XL (Podell et al., 2023). Images were generated using ComfyUI (Comfy Org, 2025) with the Euler sampler Karras et al. (2022) (30 steps), a fixed seed of 0, and a batch size of 4 to ensure reproducibility and avoid cherry-picking. These experiments demonstrate LanPaint’s practical efficacy across diverse diffusion architectures, including rectified flow models (HiDream-L1, Flux.1, Stable Diffusion 3.5) and denoising diffusion probabilistic models (Stable Diffusion XL). Additionally, we have released LanPaint as a publicly available, plug-and-play extension for ComfyUI.
@@
Figure 1 showcases the inpainting results. We also provide more examples in Appendix I. LanPaint consistently produces seamless inpainting across both DDPM-based (Stable Diffusion XL) and rectified flow-based (Stable Diffusion 3.5, Flux.1, HiDream-L1) architectures, highlighting its robust generalization capabilities.
5 Related Works
5.1 ODE-based Sampling Methods and Recitified Flow
Vanilla DDPMs are slow, often requiring hundreds of denoising steps. To address this, acceleration strategies have emerged, such as approximate diffusion processes (Song et al., 2020b; Liu et al., 2022a; Song et al., 2020a; Zhao et al., 2023) and advanced ODE solvers (Karras et al., 2022; Lu et al., 2022; Zhao et al., 2023). These methods convert stochastic DDPM sampling into deterministic ODE flows, enabling larger time steps and faster generation, and currently serve as the predominant sampling methods for diffusion models.
Rectified flow (Lipman et al., 2022; Liu et al., 2022b) has recently emerged as an alternative framework for diffusion models. However, Gao et al. (2025) demonstrated that rectified flows are essentially reparameterizations of existing ODE-based diffusion models, offering improved numerical properties. Thus, they can be considered part of the same family as ODE sampling methods.
LMC
One family of work tackling this problem is Langevin dynamics Monte Carlo (LMC). This approach was pioneered by RePaint (Lugmayr et al., 2022), which employs a "time travel" mechanism of iterative denoising and renoising steps. This mechanism is later known (Cornwall et al., 2024) to be equivalent to LMC. A crucial advantage of this formulation is that it enables easy switching between stochastic and ODE sampling, by simply switching the denoising step from SDE to ODE.
In this work, we first adapt both RePaint and the Langevin dynamics approach from Cornwall et al. (2024) to an ODE sampler as our baseline. Through this implementation, we identify their common limitation—susceptibility to local likelihood maxima trapping—and subsequently address it with our proposed bidirectional guidance.
SMC
Linear Inverse Problems
One approach minimizes heuristic losses $\|\mathbf{y}-H\mathbf{z}\|_{2}^{2}$, as in MCG (Chung et al., 2022b), DPS (Chung et al., 2022a), GradPaint (Grechka et al., 2024), DCPS (Janati et al., 2024), and D-Flow (Ben-Hamu et al., 2024). However, most of these methods are tailored for stochastic DDPM samplers, requiring costly full-model differentiation or expensive optimization (e.g., line search). For our baselines, we select only those compatible with ODE samplers and free from line search.
Another approach uses variational inference in the DDIM framework, such as DDRM (Kawar et al., 2022). CoPaint (Zhang et al., 2023a) also adopt variational inference to enhance its heuristic losses optimization. While DDIM enables fast deterministic sampling, its variational approximations introduce limitations.
While we include these inverse problem baselines for comparison, they differ fundamentally from partial conditional sampling. Linear inverse problems (effective for inpainting) do not enforce matching the joint distribution between inpainted and known regions—a core requirement of our approach. Accordingly, these baselines are included as supplementary reference points rather than essential benchmarks.
5.3 Trained Partial Conditional Sampling with Diffusion Models
Limitation
@@ 5.3 Trained Partial Conditional Sampling with Diffusion Models
LanPaint’s exactness comes at a cost: it heavily relies on the score interpretation of diffusion models. This interpretation, while applicable to various architectures such as variance-preserving, variance-exploding, and flow matching, is valid only for models trained from scratch, not for distilled models trained without denoising or flow matching. Our experience shows that LanPaint’s performance degrades with distilled models.
Appendix A Diffusion Process and Langevin Dynamics
@@
This is derived by differentiating $f$ using the chain rule with the help of equation 16 and equation 18, while keeping all terms up to order $dt$ (note that $d\mathbf{W}$ scales as $\sqrt{dt}$). The emergence of the second-order derivative term $\nabla_{\mathbf{x}}^{2}f$ is the key distinction from ordinary calculus. We will later use this lemma to analyze the evolution of the distribution of $\mathbf{x}_{t}$.
@@
To verify stationarity, we will show that after evolution from time 00 to $\Delta t$, the distribution of $\mathbf{x}_{\Delta t}$ is still $p(\mathbf{x})$. Consider a test function $f$ and initial positions $\mathbf{x}_{0}\sim p(\mathbf{x})$, stationary can be assessed by tracking the change in the expectation $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x})}[f(\mathbf{x}_{\Delta t})]$. Using Itô’s lemma and note that $\mathbb{E}_{\mathbf{x}}[d\mathbf{W}]=\mathbf{0}$ for any distribution of $\mathbf{x}$, we compute:
[…]
where 00 is obtained by substituting $\mathbf{s}=\nabla_{\mathbf{x}}\log p$. Because $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x})}\left[f(\mathbf{x}_{\Delta t})-f(\mathbf{x}_{0})\right]=0$ for any test function $f$, this means the distribution of $\mathbf{x}_{\Delta t}$ must have been kept the same as $\mathbf{x}_{0}$.
@@
Appendix B The Denoising Diffusion Probabilistic Model (DDPMs)

### v3

@@ 5.3 Latent and Pixel Space Model: CelebA and ImageNet
Table 2: LPIPS and FID comparison on ImageNet for various inpainting and outpainting setups. Euler Discrete Sampler, 20 steps. Lower LPIPS and FID values indicate better perceptual similarity and feature distribution similarity to the ground truth, respectively. Numerical suffixes (-5, -10) denote inner iterations (network evaluations per sampling step, except for CoPaint, which requires multiple evaluations per inner iteration). Time per sample and memory overhead (extra memory required during inference) are also reported. Evaluations were conducted on a single RTX 3090.
@@
LanPaint-5 (ours) 0.180 49.3 0.323 49.3 0.127 24.1 0.508 68.6 11.3 599
LanPaint-10 (ours) 0.171 46.4 0.314 44.7 0.117 21.2 0.486 62.0 20.8 599
We evaluate the inpainting performance of LanPaint on the CelebA-HQ-256 (Liu et al., 2015) and ImageNet-256 (Deng et al., 2009) datasets, leveraging pre-trained latent (Rombach et al., 2021) and pixel space (Dhariwal & Nichol, 2021) diffusion models, respectively. The experiments assess reconstruction quality across various mask geometries, including box, half, checkerboard, and outpainting. Following the same setting as the previous works (Kawar et al., 2022; Chung et al., 2022b), perceptual fidelity is quantified through LPIPS (Zhang et al., 2018) and FID metrics, calculated on 1,000 validation images per dataset. Results are presented in Tables 1 and 2. We also provide qualitative visualization of generated samples in Fig.5 and Fig.6. All methods employ consistent parameters across tasks and masks, utilizing a 20-step Euler Discrete Sampler. Further details about parameters are provided in Appendix A.
LanPaint consistently achieves superior LPIPS and FID scores across most test scenarios, demonstrating robustness, particularly in challenging checkerboard and outpainting tasks. In contrast, methods such as DPS (Chung et al., 2022a) and CoPaint (Zhang et al., 2023a), designed for stochastic sampling with 250–500 steps, exhibit reduced performance in the 20-step ODE setting.
Notably, DPS and CoPaint’s removing of the manifold constraint from MCG (Chung et al., 2022b) compromises their stability, leading to poorer performance compared to MCG, which remains robust among heuristic methods. This contradicts prior findings in stochastic DDPM sampling, where removing the manifold constraint typically enhances performance.
On CelebA-HQ-256, Replace (Song & Ermon, 2019b) marginally outperforms LanPaint in FID for the half-mask scenario (30.3 vs. 30.4), a result attributed to FID’s high variance when large inpainted regions deviate significantly from the original images. This variance is exacerbated by the use of 1,000 validation images, as opposed to the typical 30,000 for highly divergent sets, similarly affecting outpainting results. Consequently, FID scores for half and outpainting masks should be interpreted cautiously.
@@ 5.4 Ablation Study
Table 3 presents the ablation study of LanPaint’s major components: the BiG score and FLD, along with the impact of step size on performance. The study is conducted on two datasets, CelebA-HQ-256 and ImageNet, using the box inpainting tasks (Other masks share similar trends). Results are reported for five different step sizes (0.02, 0.05, 0.1, 0.15, and 0.2) with LPIPS and FID metrics, where lower values indicate better performance. Sensitivity of other parameters is provided in Fig.8.
For the ImageNet dataset, the ablation study demonstrates that both the BiG score and FLD significantly contribute to overall performance. Without FLD, the Langevin dynamics diverge as the step size increases from 0.05 to 0.15, resulting in progressively worse performance metrics. However, incorporating FLD suppresses this divergence, enabling the use of larger step sizes and improving performance. The BiG score enhances performance by facilitating bidirectional information flow between the inpainted and known regions, while FLD supports larger step sizes, accelerating the convergence of the Langevin dynamics and yielding better results with the same number of iterations.
@@
Table 3: Ablation study of LanPaint-10’s components on CelebA-HQ-256 and ImageNet with box inpainting. Results for different step sizes (0.02, 0.05, 0.1, 0.15, 0.2) are shown with LPIPS and FID metrics. Lower values are better.
Step Size 0.02 Step Size 0.05 Step Size 0.1 Step Size 0.15 Step Size 0.2
Method LPIPS FID LPIPS FID LPIPS FID LPIPS FID LPIPS FID
@@
None (Langevin) 0.121 28.9 0.115 28.4 0.111 29.2 0.114 30.9 0.108 30.1
+ BiG score 0.110 26.1 0.104 26.5 0.102 28.5 0.103 28.7 0.103 30.0
+ FLD 0.121 28.6 0.115 28.9 0.112 29.9 0.116 31.7 0.109 30.5
+ (BiG score + FLD) 0.111 26.7 0.105 26.6 0.103 28.5 0.103 29.5 0.103 30.2
@@
None (Langevin) 0.220 66.7 0.223 65.1 0.314 81.6 0.441 125.9 0.475 141.1
+ BiG score 0.205 58.6 0.213 58.0 0.303 76.0 0.431 121.4 0.474 140.0
+ FLD 0.217 68.3 0.205 60.7 0.195 56.9 0.188 54.4 0.181 51.4
+ (BiG score + FLD) 0.201 58.9 0.190 52.7 0.179 48.1 0.171 46.4 0.167 45.1
@@ 5.5 Production-Level Model Evaluation Across Architectures: Stable Diffusion, Flux, and HiDream
Previous evaluations of LanPaint were limited to academic benchmarks, leaving its performance on real-world production diffusion models—with their diverse architectures and higher resolutions—unexamined. Notably, to the best of our knowledge, no prior training-free inpainting methods, except variants of the replace method (built in ComfyUI), have demonstrated such validation in their publications or been implemented by third parties for this purpose.
@@
To assess LanPaint’s effectiveness on modern generative models, we implemented it on HiDream-L1 (HiDream.ai, 2025), Flux.1 Dev (Labs, 2024), Stable Diffusion 3.5 (Esser et al., 2024), and Stable Diffusion XL (Podell et al., 2023). Images were generated using ComfyUI (Comfy Org, 2025) with the Euler sampler (Karras et al., 2022) (30 steps), a fixed seed of 0, and a batch size of 4 to ensure reproducibility and avoid cherry-picking. These experiments demonstrate LanPaint’s practical efficacy across diverse diffusion architectures, including rectified flow models (HiDream-L1, Flux.1, Stable Diffusion 3.5) and denoising diffusion probabilistic models (Stable Diffusion XL). Additionally, we have released LanPaint as a publicly available, plug-and-play extension for ComfyUI.
@@
Fig.1 showcases the inpainting results. We also provide more examples in Appendix I. LanPaint consistently produces seamless inpainting across both DDPM-based (Stable Diffusion XL) and rectified flow-based (Stable Diffusion 3.5, Flux.1, HiDream-L1) architectures, highlighting its robust generalization capabilities.
Figure 7: LanPaint-2 video inpainting and outpainting on Wan 2.2 T2V 14B models (seed=0, 480p resolution, 20 diffusion sampling steps). Top: Inpainting with prompt "Add a white fascinator" showing input with masked region (cyan) and output. Bottom: Outpainting with prompt "Change ratio to 11:6" showing input with padding regions (cyan) and output. Five keyframes displayed (frames 1, 10, 20, 30, 40).
5.6 Video Inpainting Results
We demonstrate LanPaint’s dimension-agnostic formulation on video inpainting and outpainting using the Wan 2.2 T2V model  Wan et al. (2025) (14B parameters, fp8-scaled). Our setup processes 40-frame, 480p clips using Euler ODE sampling with 20 steps and two LanPaint inner iterations. Fig. 7 highlights these capabilities. These results affirm LanPaint’s versatility for video editing. Full videos are available at https://github.com/scraed/LanPaint/tree/master/examples.
Limitation and Future work
@@ 5.6 Video Inpainting Results
LanPaint’s exactness comes at a cost: it heavily relies on the score interpretation of diffusion models. This interpretation, while applicable to various architectures such as variance-preserving, variance-exploding, and flow matching, is valid only for models trained from scratch, not for distilled models trained without denoising or flow matching. Our experience shows that LanPaint’s performance degrades with distilled models. Future studies on distillation methods that preserve the score interpretation of diffusion models are desired. A distillation method that captures LanPaint’s capabilities within a model is also of interest, as it could significantly accelerate LanPaint.
LanPaint assumes noise-free observations $\mathbf{y}_{o}$. Adapting it to handle noisy observations with a specified noise level is feasible but requires modifying the conditional distribution $p_{t}(\mathbf{y}\mid\mathbf{y}_{o})$ in Eq.11. This adaptation represents a promising direction for future research.
In this paper, we primarily focus on image inpainting. However, LanPaint, as a conditional sampling method independent of data modality, can be applied to diverse domains, including text, audio, video, and scientific applications such as protein scaffolding and fluid field reconstruction.
Broader Impact Statement
LanPaint’s efficient image inpainting boosts creative applications but risks misuse in generating deepfakes or misinformation. We advocate watermarking, provenance tracking, and community regulation to mitigate harm, as discussed in (Denton, 2021) and (Franks & Waldman, 2018).
Appendix A More Ablation and Implementation Details
Mask Types
LanPaint
We implement LanPaint using the diffusers package, following Algorithm 4. Hyperparameters are configured as follows: $\gamma=15$, $\alpha=0.$, and $\lambda=8$ for all image inpainting tasks, drawing loosely from the insights gained through sensitivity analysis in Fig.8. The notation LanPaint-5 and LanPaint-10 denotes $N=5$ and $N=10$ sampling steps, respectively. The step size $\eta$ is set to $0.15$ for both Celeb-A and ImageNet. The impact of step size is ablated in Table 3, with the impact of other parameters discussed in Fig.8. We have also provided impact of different samplers in Table.4.
Figure 8: Impact of expected noise $\alpha$, guidance scale $\lambda$, and friction $\gamma$ on LanPaint-10’s LPIPS for ImageNet box inpainting at stepsize 0.15, evaluated on a validation set of 100 images. Expected noise $\alpha$ most affects LPIPS, ideally 0 for image inpainting. Guidance scale $\lambda$ significantly improves performance from 0 (no guidance) to 4, with an optimal range of 6–10. The friction parameter $\gamma$ has a less significant effect; in practice, we use a prescribed value of $\gamma=15$ without finetuning, sharing this value across all tasks.
Table 4: Ablation study on the impact of different diffusion samplers (Euler, DPM++Karras(Lu et al., 2025), and DDIM (Song et al., 2020b)) for various heuristic and asymptotically exact inpainting methods. Performance is evaluated using LPIPS (lower is better) and FID (lower is better) metrics on 1,000 images from the CelebA and ImageNet datasets with box mask.
CelebA ImageNet
Euler DPM++ DDIM Euler DPM++ DDIM
Method LPIPS FID LPIPS FID LPIPS FID LPIPS FID LPIPS FID LPIPS FID
Heuristic Methods
Replace 0.131 31.7 0.119 26.7 0.130 31.5 0.229 75.7 0.234 75.1 0.228 75.6
CoPaint-2 0.180 43.6 0.186 45.8 0.169 41.1 0.234 85.1 0.233 84.5 0.297 89.1
CoPaint-3 0.172 41.7 0.181 44.3 0.163 38.9 0.228 76.9 0.228 78.9 0.288 82.1
DDRM 0.128 32.4 0.135 34.8 0.127 32.2 0.216 67.2 0.215 68.9 0.211 65.3
MCG 0.130 31.6 0.113 25.9 0.124 29.7 0.225 83.5 0.266 88.1 0.253 86.7
DPS 0.181 44.1 0.166 40.2 0.179 43.7 0.252 107.5 0.248 103.0 0.249 108.2
Asymptotically Exact Methods
Repaint-5 0.115 34.5 0.127 41.7 0.114 34.2 0.216 62.8 0.342 133.6 0.211 60.8
Repaint-10 0.112 34.6 0.409 223.3 0.111 34.4 0.215 61.0 0.604 174.5 0.211 59.7
TFG-5 0.119 31.9 0.113 31.9 0.118 31.4 0.235 69.3 0.281 83.7 0.234 69.7
TFG-10 0.114 33.2 0.108 33.2 0.112 31.2 0.234 66.5 0.305 89.5 0.235 65.8
LanPaint-5 (ours) 0.105 27.9 0.097 23.7 0.104 28.0 0.180 49.3 0.201 56.2 0.178 48.6
LanPaint-10 (ours) 0.103 29.5 0.097 26.9 0.103 29.3 0.171 46.4 0.205 55.1 0.172 45.6
TFG
TFG ((Cornwall et al., 2024)) corresponds to pure Langevin sampling. It is implemented via the Euler-Maruyama discretization, with step size schedule $\mathrm{d}\tau=\eta\sqrt{1-\bar{\alpha}_{t}}$ with $\eta=0.04$ as reported in their paper.
DDRM
Our implementation of DDRM follows Equations (7) and (8) in Kawar et al. (Kawar et al., 2022), with $\sigma_{y}=0$, $\mathbf{V}=I$, and $s_{i}=1$ for observed regions $\mathbf{y}$ and $s_{i}=0$ for inpainted regions $\mathbf{x}$. The hyperparameters $\eta=0.7$ and $\eta_{b}=1$ are selected based on the optimal KID scores reported in Table 3 of the reference.
MCG
We implement MCG as described in Algorithm 1 of Chung et al. (Chung et al., 2022b), using $\alpha=0.1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$ as recommended in Appendix C. An alternative choice, $\alpha=0.1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$, was evaluated but resulted in significantly poorer performance on ImageNet, as shown in Table 5. Consequently, we opted against using this alternative.
DPS
We adopt DPS (Chung et al., 2022a) using $\alpha=1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$ as recommended in Appendix D.
Alpha 0.1 Alpha 1.0
Method LPIPS FID LPIPS FID
CelebA-HQ-256
MCG 0.130 31.6 0.125 30.1
DPS 0.190 41.5 0.181 44.1
ImageNet
MCG 0.226 83.5 0.240 79.5
DPS 0.251 110.7 0.252 107.5
RePaint
We implement RePaint based on Algorithm 1 in Lugmayr et al. (Lugmayr et al., 2022), with modifications to accommodate fast samplers. While the original method was designed for DDPM, we adapt it by replacing the backward sampling step (Line 7) with a Euler Discrete Sample step. Additionally, we set the jump step size to $1$ instead of the default $10$ backward steps recommended in Appendix B of the original work. This adjustment is necessary because fast samplers typically operate with only around $20$ backward sampling steps, making larger jump sizes impractical.
Appendix B Diffusion Process and Langevin Dynamics
@@
This is derived by differentiating $f$ using the chain rule with the help of Eq.18 and Eq.20, while keeping all terms up to order $dt$ (note that $d\mathbf{W}$ scales as $\sqrt{dt}$). The emergence of the second-order derivative term $\nabla_{\mathbf{x}}^{2}f$ is the key distinction from ordinary calculus. We will later use this lemma to analyze the evolution of the distribution of $\mathbf{x}_{t}$.
@@
To verify stationarity, we will show that after evolution from time 0 to $\Delta t$, the distribution of $\mathbf{x}_{\Delta t}$ is still $p(\mathbf{x})$. Consider a test function $f$ and initial positions $\mathbf{x}_{0}\sim p(\mathbf{x})$, stationary can be assessed by tracking the change in the expectation $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x})}[f(\mathbf{x}_{\Delta t})]$. Using Itô’s lemma and note that $\mathbb{E}_{\mathbf{x}}[d\mathbf{W}]=\mathbf{0}$ for any distribution of $\mathbf{x}$, we compute:
$\begin{split}\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x})}\left[f(\mathbf{x}_{\Delta t})-f(\mathbf{x}_{0})\right]&\approx\Delta t\int p(\mathbf{x})\left(\nabla_{\mathbf{x}}f\cdot\mathbf{s}+\nabla^{2}_{\mathbf{x}}f\right)d\mathbf{x}\\ &=\Delta t\int f(\mathbf{x})\left(-\nabla_{\mathbf{x}}\cdot(p\mathbf{s})+\nabla^{2}_{\mathbf{x}}p\right)d\mathbf{x}\quad\text{(integration by parts)}\\ &=\Delta t\int f(\mathbf{x})\nabla_{\mathbf{x}}\cdot\left(-p\mathbf{s}+\nabla_{\mathbf{x}}p\right)d\mathbf{x}\\ &=0,\end{split}$ (23)
where 0 is obtained by substituting $\mathbf{s}=\nabla_{\mathbf{x}}\log p$. Because $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x})}\left[f(\mathbf{x}_{\Delta t})-f(\mathbf{x}_{0})\right]=0$ for any test function $f$, this means the distribution of $\mathbf{x}_{\Delta t}$ must have been kept the same as $\mathbf{x}_{0}$.
@@
Appendix C The Denoising Diffusion Probabilistic Model (DDPMs)